OpenStudy (anonymous):

Write these percents as fractions 6% = 5% = 88% = 24% =

OpenStudy (anonymous):

\[\LARGE 6 \% =\frac{6}{100}\] now do the rest .. :)

OpenStudy (anonymous):

Thank you everyone! Closing.

OpenStudy (anonymous):

0.06

OpenStudy (anonymous):

he said as fractions... \[\LARGE 5\%=\frac{5}{100} \quad \quad , \quad 88\% =\frac{88}{100} \] I'll leave you to finish the last one ;)